-When a firm produces one more unit of output the total revenue increases from $803 to $1,037, and the total cost increases from $519 to $708. When this last unit of output was produced, what was the change in profit? Enter a whole number with no dollar sign. Enter a negative sign if appropriate.

-With her resources Sally can produce 36 units of good X or 142 units of good Y. What is her opportunity cost of producing 9 units of good X? Round your answer to two decimal places.

Please show steps to complete this program...
Answer #1

a) INcrease in the revenue of the firm = 1037 - 803 = 243, increase in the cost of the firm 708 - 519 = 189.

The change in the profit is 243 - 189 = $54.

b) The opportunity cost of producing 1 good X = 142 / 36 = 3.94.

So, to produce one good of X Sally has to give up 3.94 of good Y, so for 9 the opportunity cost will be 9 x 3.94 = 35.46 of good Y.